# Empowering Agentic Video Analytics Systems with Video Language Models

### v3

@@ 5.1. Tri-View Retrieval
Figure 4. An illustration of tri-view retrieval and borda counting on the retrieved events.
@@
To comprehensively retrieve relevant information from the index for a given query, Avas employs a three-view retrieval process: the first view targets events, enabling retrieval at the event level to provide information for event summary-related queries. Specifically, the query is encoded using the text encoder JinaCLIP (Koukounas et al., 2024) and matched against the events table in the constructed EKG. The second view focuses on entities, offering insights into basic facts or item-specific queries. For this, we leverage the entity centroids extracted and aggregated as detailed in §4.3 to facilitate retrieval. The third view utilizes vision embeddings of raw video frames as complementary information. The retrieved entities and raw frames are subsequently linked to their associated events through the constructed EKG.
@@
It is important to note that the retrieved events should be ranked. Ranking is not only crucial for filtering noise from the retrieved results but also essential for enabling agentic searching, as detailed in §5.2. A straightforward ranking method, such as similarity-based ranking, cannot be directly applied to Avas due to the retrieved events originating from three distinct views. To integrate these results, we propose to use a weighted Borda counting approach.
@@
Fig. 4 illustrates the process of using Borda counting to integrate and rank retrieved events from the three views in Avas. Specifically, we select the top $K$ events from each view and rank them based on their calculated similarities within that view. Subsequently, the similarities of these $K$ events are normalized to compute their Borda scores:
(2) $s_{m}(e_{j})=\frac{\text{sim}_{m}(e_{j})}{\sum_{e_{k}\in\mathcal{E}_{m}}\text{sim}_{m}(e_{k})},$
@@
Figure 5. An example of agentic tree search with four actions and a depth of three, yielding 13 distinct pathways for information gathering and response generation.
5.2. Agentic Searching on Graph
The retrieved events mentioned above can be directly utilized to generate the final answer. However, to support complex queries, such as query-focused summaries and multi-hop queries, Avas searches for additional relevant information by leveraging the relationships between events and entities within the constructed EKG. To enable efficient exploration, we propose the agentic searching on graph approach.
This approach is inspired by human behavior when retrieving and reasoning about information in videos. Specifically, individuals typically begin by identifying key clips based on retrieval keywords, then gather additional context by exploring clips preceding or following the identified ones. To achieve a more comprehensive understanding, they may also refine their search by using alternative keywords, iterating through this process as needed. Similarly, in our agentic searching process, we define the agentic action space as follows:
@@ 5.2. Agentic Searching on Graph
Re-query (RQ): this action generates a new query represented by a list of keywords via a LLM and retrieves complementary events as outlined in §5.1. It reflects the human tendency to gather information from multiple perspectives to achieve a more comprehensive understanding.
Summary and answer (SA): this action utilizes the descriptions of the retrieved events from the EKG and generates the response to the specific query by employing a LLM.
By leveraging these predefined agentic actions, we design the agentic searching process as a tree search framework. The search begins with an initial retrieval based on the original query, producing a list of relevant events that serves as the root node of the search tree. At each depth level, a one-step rollout of all four predefined actions is performed for each active node. If the SA action is reached, the corresponding search path is terminated. The rollout process continues until the maximum tree depth is reached. This tree search process explores diverse pathways to gather information from the EKG; As iilustrated in Fig. 5, a tree with a depth of 3 would result in 13 distinct information-gathering paths and corresponding answers. All generated answers are evaluated, and the optimal one is selected using the thought consistency method, which is detailed in §5.3.
@@
5.3. Consistency Enhanced Generation
@@
At each SA node, instead of generating the answer a single time, we repeatedly generate answers multiple times using a Chain-of-Thought (CoT) prompting scheme. Following the principle of self-consistency (Wang et al., 2023), correct answers are more likely to emerge consistently across multiple valid reasoning trajectories during repeated generations. Specifically, we evaluate the consistency not only across the generated answers but also within their associated CoT traces. Nodes demonstrating strong internal coherence, where the reasoning process aligns logically with the conclusion, are assigned higher scores.
@@
(5) $S_{r}^{(t)}=\frac{2}{k(k-1)}\sum\limits_{1\leq i<j\leq k}\text{BERTScore}(r_{i},r_{j}),$
@@
(6) $S_{\text{final}}^{(t)}=\lambda S_{a}^{(t)}+(1-\lambda)S_{r}^{(t)},$
where $\lambda\in[0,1]$ is a weighting parameter controlling the trade-off between answer agreement and thought consistency. In our implementation, we set it to 0.3, the parameter tuning would be discussed in §7.4.3.
@@
6. Implementation
Avas uses Qwen2.5-VL-7B for constructing EKGs, Qwen2.5-32B for SA, and Gemini-1.5-Pro for CA. Local VLMs and LLMs are deployed using LMDeploy(Contributors, 2023) to enable accelerated inference. Additionally, Avas adopts batch inference for several key stages—including description generation, description merging, entity extraction, and tree search—to improve efficiency and maximize GPU utilization. For text and vision embedding, we utilize JinaCLIP(Koukounas et al., 2024). And we employ BERTScore with the deberta-xlarge-mnli(He et al., 2020) checkpoint. The storage of EKG and vector representations is based on the implementation of (Guo et al., 2024), upon which we make further modifications to suit the specific requirements of Avas.
@@ 6. Implementation
”…Your task is to extract and provide a detailed description of the video segment, focusing on all key visible details…”.
@@
For scenario-specific videos, we design the prompts to emphasize scenario-relevant information. For example, in the case of wildlife surveillance scenario, key information may includes the timestamps of the recording, animal activities (e.g., presence, species, number, specific behaviors, etc.), and environmental changes.
7. Evaluation
7.1. Evaluation Settings
@@
Figure 6. The achieved accuracy of Avas and various baselines on the LVBench, VideoMME-Long, and Avas-100 benchmarks.
7.1.1. Benchmark
@@
Avas is evaluated on two widely used public long-video benchmark and one ultra-long video benchmarks proposed by us, covering a broad range of video scenarios and problem types.
@@
LVBench(Wang et al., 2024b) stands out among publicly available benchmarks for its exceptionally long average video duration, approximately 4100 seconds per video. It comprises 103 videos with a total of 1549 questions, covering six distinct video domains and addressing six task types including temporal grounding, summarization, and reasoning.
@@
VideoMME-Long(Fu et al., 2024) is a subset of the VideoMME benchmark, focusing on videos exceeding 20 minutes in duration, with an average length of 2400 seconds. Comprising a total of 300 videos and 900 questions, the benchmark covers a wide range of video themes, spanning 6 primary visual domains with 30 subfields to ensure broad scenario generalizability, and includes 12 distinct task types.
@@
Avas-100 is proposed by us, which is an ultra-long video benchmark specially designed to evaluate video analysis capabilities Avas-100 consists of 8 videos, each exceeding 10 hours in length, and includes a total of 120 manually annotated questions. The benchmark covers four typical video analytics scenarios: human daily activities, city walking, wildlife surveillance, and traffic monitoring, each scenario contains two videos. The human daily activity scenario features egocentric videos selected and stitched from the Ego4D(Grauman et al., 2022). City walking and wildlife surveillance videos are curated from publicly available recordings on YouTube, capturing urban exploration and animal monitoring respectively. Traffic monitor videos are composed from clips in the Bellevue Traffic Video Dataset (Bhardwaj et al., 2022). All questions are carefully designed by human annotators, who also provide reference answers as the ground truth. In addition, GPT-4o is utilized to generate plausible distractor options. The accuracy is evaluated by analyzing Avas’s responses to multiple-choice questions included in the benchmarks.
7.2. Baselines
@@
We conduct a comprehensive comparison between Avas and a wide range of baseline models, encompassing both mainstream VLMs and specialized Video-RAG methods. The VLM baselines include GPT-4o (Achiam et al., 2023), Gemini-1.5-Pro (Team et al., 2023), Phi-4-Multimodal (Abdin et al., 2024b), Qwen2.5-VL-7B (Bai et al., 2025), InternVL2.5-8B (Chen et al., 2024), and LLaVA-Video-7B (Zhang et al., 2024a). Each of these models is evaluated with two typical strategies: uniform sampling and vectorized retrieval, where a CLIP-based retriever selects the top-K relevant frames based on the user query. In addition to VLMs, we benchmark Avas against SOTA Video-RAG frameworks, including VideoTree(Wang et al., 2024c), VideoAgent(Wang et al., 2024d), DrVideo(Ma et al., 2024), and VCA(Yang et al., 2024). Among these, VideoTree, VideoAgent, and VCA are built upon GPT-4o, while DrVideo leverages GPT-4.
7.3. Overall Evaluation
7.3.1. Overall Performance
Fig. 6 illustrates the overall accuracy achieved by Avas compared to various baselines on the LVBench, VideoMME-Long, and Avas-100 benchmarks. Across all three benchmarks, Avas consistently outperforms the baselines. Specifically, on LVBench, Avas delivers a remarkable 16.9% improvement, while on VideoMME-Long, it advances the SOTA by approximately 5.2%. On the Avas-100 benchmark, Avas achieves an accuracy of 75.8%, significantly surpassing all competing methods.
In detail, compared to video-RAG methods, Avas achieves improvements of 21% and 7.8% on LVBench and VideoMME-Long, respectively. When compared to vectorized retrieval-based methods, Avas demonstrates gains of 16.9% on LVBench and 20.8% on Avas-100. Furthermore, against uniform sampling baselines, Avas improves performance by approximately 19.6% and 26.9% on LVBench and Avas-100, respectively.
Notably, on Avas-100, when evaluated with extremely long videos, Avas maintains robust performance, whereas the baselines degrade significantly. This highlights the effectiveness of Avas in handling L4 video analytics tasks.
@@ 7.3.1. Overall Performance
Figure 7. The accuracy achieved by Avas and the baselines across typical query categories on LVBench: Temporal Grounding (TG), Summarization (SU), Reasoning (RE), Entity Recognition (ER), Event Understanding (EU), and Key Information Retrieval (KIR).
7.3.2. Performance on Different Query Categories
@@
We also evaluate the accuracy achieved by Avas across typical query categories on LVBench. As illustrated in Fig. 7, our approach achieves improvements of 16%, 5.3%, 35.6%, 21.2%, 17.5%, and 18.9% across six key task types: Temporal Grounding, Summarization, Reasoning, Entity Recognition, Event Understanding, and Key Information Retrieval, respectively, compared to the uniform sampling and vectorized retrieval baselines powered by Gemini-1.5-Pro. Notably, Avas demonstrates particularly strong performance on reasoning tasks, which require identifying causal relationships between events and linking preceding and succeeding events within the video. This highlights Avas’s ability to effectively locate and extract critical information from long videos, thereby enabling advanced L4 video analytics systems.
@@
Figure 8. The accuracy achieved by Avas and baselines across three benchmarks when utilizing different LLMs and VLMs.
7.3.3. Performance under Different Configurations
@@
Fig. 8 shows the performance of Avas using different models configurations for SA and CA. For SA, two models were used: Qwen2.5 14B and 32B. For CA, two models were used: Qwen2.5-VL-7B and Gemini-1.5-Pro. The results show that across the three benchmarks, Avas using Gemini-1.5-Pro for CA achieved improvements of 18.9%, 5.2%, and 20.8% respectively compared to the best baseline result using the same model, while using Qwen2.5-VL-7B yielded improvements of 13%, 7.2%, and 15% respectively, fully demonstrating the effectiveness of our method. Notably, even when only using Qwen2.5-32B and Qwen2.5-7B based on the textual content from EKG without accessing raw frames, Avas can surpass the performance of Qwen2.5-VL-7B on the three benchmarks and also outperform most models shown in Figs. 6(a), 6(b), and 6(c).
7.3.4. Performance on Different Video Lengths
@@
Figure 9. The accuracy achieved by Avas and the baselines across varying video lengths via concatenating videos from LVBench.
@@
To evaluate the robustness of Avas with respect to video length, we conducted experiments on videos of varying durations. Specifically, sequences of 3.3, 6.6, and 10 hours were created by concatenating videos from the VideoMME-Long benchmark. Performance was measured using identical questions across these varying video lengths. As illustrated in Fig. 9, both Qwen2.5-VL-7B and Gemini-1.5-Pro baselines exhibit significant performance degradation as video length increases. When extended to 10 hours, their performance declines by 4.6% and 8.2%, respectively, under the uniform sampling method, compared to the original VideoMME-Long benchmark. For the vectorized retrieval setting, the performance drops are 4.6% and 5.5%, respectively. These results highlight the limitations of these methods in scaling effectively with increasing video length. In contrast, Avas consistently maintains stable performance across all video lengths, underscoring its robustness and scalability in handling video data of any duration.
7.3.5. System Overhead

### v5

@@ 5.1 Tri-View Retrieval
Figure 5: An illustration of tri-view retrieval and borda counting on the retrieved events.
@@
To comprehensively retrieve relevant information from the index for a given query, Ava employs a tri-view retrieval process: the first view targets events, enabling retrieval at the event level to provide information for event summary-related queries. Specifically, the query is encoded using the text encoder JinaCLIP [22] and matched against the events table in the constructed EKG. The second view focuses on entities, offering insights into basic facts or item-specific queries. For this, we leverage the entity centroids extracted and aggregated as detailed in §4.3 to facilitate retrieval. The third view utilizes vision embeddings of raw video frames as complementary information. The retrieved entities and raw frames are subsequently linked to their associated events through the constructed EKG.
@@
It is important to note that the retrieved events should be ranked. Ranking is not only crucial for filtering noise from the retrieved results but also essential for enabling agentic searching, as detailed in §5.2. A straightforward ranking method, such as similarity-based ranking, cannot be directly applied to Ava due to the retrieved events originating from three distinct views. To integrate these results, we propose to use a weighted Borda counting approach.
@@
Fig. 5 illustrates the process of using Borda counting to integrate and rank retrieved events from the three views in Ava. Specifically, we select the top $K$ events from each view and rank them based on their calculated similarities within that view. Subsequently, the similarities of these $K$ events are normalized to compute their Borda scores:
$s_{m}(e_{j})=\frac{\text{sim}_{m}(e_{j})}{\sum_{e_{k}\in\mathcal{E}_{m}}\text{sim}_{m}(e_{k})},$ (2)
@@
Figure 6: An example of agentic tree search with four actions and a depth of three, yielding 13 distinct pathways for information gathering and response generation.
5.2 Agentic Searching on Graph
The retrieved events mentioned above can be directly utilized to generate the final answer. However, to support complex queries, such as query-focused summaries and multi-hop queries, Ava searches for additional relevant information by leveraging the relationships between events and entities within the constructed EKG. To enable efficient exploration, we propose the agentic searching on graph approach.
The proposed approach is inspired from human strategies for information retrieval and reasoning within video content. Typically, individuals initiate this process by locating salient segments guided by retrieval keywords (i.e., query), subsequently enriching their understanding by examining the temporal context in adjacent clips (i.e., forward and backward). For a more thorough comprehension, they may iteratively refine their search using alternative keywords as necessary (i.e., re-query). Similarly, in our agentic searching process, we define the agentic action space as follows:
@@ 5.2 Agentic Searching on Graph
Re-query (RQ): this action generates a new query represented by a list of keywords via an LLM and retrieves complementary events as outlined in §5.1. It reflects the human tendency to gather information from multiple perspectives to achieve a more comprehensive understanding.
Summary and Answer (SA): this action utilizes the descriptions of the retrieved events from the EKG and generates the response to the specific query by employing an LLM.
Utilizing these predefined agentic actions, we structure the agentic search process within a tree search framework. The search is initiated by performing an initial retrieval based on the original query, yielding a set of relevant events that constitute the root node of the search tree, as outlined in §5.1. For each node, a single-step rollout is conducted in which four predefined actions, namely, forward, backward, re-query, and summary and answer, are executed on every events on the node. Upon reaching the SA action, the corresponding search trajectory is terminated by generating an answer. The rollout proceeds iteratively until the predefined maximum tree depth is reached. This tree search mechanism systematically explores multiple pathways to extract information from the EKG. As depicted in Fig. 6, a tree of depth 3 yields 13 distinct information-gathering paths along with their respective answers. All generated answers are subsequently evaluated, with the optimal answer selected via the thought consistency method elaborated in §5.3.
@@
5.3 Consistency Enhanced Generation
@@
At each SA node, instead of generating the answer a single time, we repeatedly generate answers multiple times using a Chain-of-Thought (CoT) prompting scheme. Following the principle of self-consistency [46], correct answers are more likely to emerge consistently across multiple valid reasoning trajectories during repeated generations. Specifically, we evaluate the consistency not only across the generated answers but also within their associated CoT traces. Nodes demonstrating strong internal coherence, where the reasoning process aligns logically with the conclusion, are assigned higher scores.
@@
$S_{r}^{(t)}=\frac{2}{k(k-1)}\sum\limits_{1\leq i<j\leq k}\text{BERTScore}(r_{i},r_{j}),$ (5)
@@
$S_{\text{final}}^{(t)}=\lambda S_{a}^{(t)}+(1-\lambda)S_{r}^{(t)},$ (6)
where $\lambda\in[0,1]$ is a weighting parameter controlling the trade-off between answer agreement and thought consistency. In our implementation, we set it to 0.3, the parameter tuning would be discussed in §8.
@@
6 Implementation
In Ava we use Qwen2.5-VL-7B for constructing EKGs, Qwen2.5-32B for SA, and Gemini-1.5-Pro for CA. We utilize AWQ [24] and LMDeploy [10] to accelerate the on-device inference. These particular LLMs are selected for Ava due to our two-step principle: we first determine an appropriate model size that is feasible for edge deployment, and then, within that size range, choose the best-performing model available according to the public benchmarks [14, 35]. Additionally, Ava adopts batch inference for several key stages—including description generation, description merging, entity extraction, and tree search—to improve efficiency and maximize GPU utilization. For text and vision embedding, we utilize JinaCLIP[22]. And we employ BERTScore with the deberta-xlarge-mnli[18] checkpoint. The storage of EKG and vector representations is based on the implementation of [16], upon which we make further modifications to suit the specific requirements of Ava.
@@ 6 Implementation
"…Your task is to extract and provide a detailed description of the video segment, focusing on all key visible details…".
@@
For scenario-specific videos, we design the prompts to emphasize scenario-relevant information. For example, in the case of wildlife Monitoring scenario, key information may includes the timestamps of the recording, animal activities (e.g., presence, species, number, specific behaviors, etc.), and environmental changes. Particularly the prompts used in Ava are listed in §A.3
The hyperparameters in Ava are determined through a parameter tuning procedure conducted on the selected datasets. Specifically, for the semantic chunking methodology described in §4.2, we set the BERTScore threshold for merging and summarizing descriptions at 0.65. Empirically, this threshold yielded partitions that exhibited greater alignment with human annotations across various scenarios. Additionally, the maximum tree search depth is set to 3, more details are provided in §7.4.2.
7 Evaluation
7.1 Evaluation Settings
(a) LVBench
(b) VideoMME-Long
(c) Ava-100
@@ 7.1 Evaluation Settings
Figure 7: The achieved accuracy of Ava and various baselines on the LVBench, VideoMME-Long, and Ava-100 benchmarks.
7.1.1 Benchmarks
@@
Ava is evaluated on two widely used public long-video benchmarks and one ultra-long video benchmarks proposed by us, covering a broad range of video scenarios and problem types.
@@
LVBench[43] stands out among publicly available benchmarks for its exceptionally long average video duration, approximately 4100 seconds per video. It comprises 103 videos with a total of 1549 questions, covering six distinct video domains and addressing six task types including temporal grounding, summarization, and reasoning.
@@
VideoMME-Long[14] is a subset of the VideoMME benchmark, focusing on videos exceeding 20 minutes in duration, with an average length of 2400 seconds. Comprising a total of 300 videos and 900 questions, the benchmark covers a wide range of video themes, spanning 6 primary visual domains with 30 subfields to ensure broad scenario generalizability, and includes 12 distinct task types.
@@
Ava-100 is proposed by us, which is an ultra-long video benchmark specially designed to evaluate video analysis capabilities Ava-100 consists of 8 videos, each exceeding 10 hours in length, and includes a total of 120 manually annotated questions. The benchmark covers four typical video analytics scenarios: human daily activities, city walking, wildlife monitoring, and traffic monitoring, each scenario contains two videos. The human daily activity scenario features egocentric videos selected and stitched from the Ego4D[15]. City walking and wildlife monitoring videos are curated from publicly available recordings on YouTube, capturing urban exploration and animal monitoring respectively. Traffic monitor videos are composed from clips in the Bellevue Traffic Video Dataset [7]. All questions are carefully designed by human annotators, who also provide reference answers as the ground truth. In addition, GPT-4o is utilized to generate plausible distractor options. The accuracy is evaluated by analyzing Ava’s responses to multiple-choice questions included in the benchmarks. More details regarding Ava-100 are provided in §A.
7.2 Baselines
@@
We conduct a comprehensive comparison between Ava and a wide range of baseline models, encompassing both mainstream VLMs and specialized Video-RAG methods. The VLM baselines include GPT-4o [5], Gemini-1.5-Pro [39], Phi-4-Multimodal [3], Qwen2.5-VL-7B [6], InternVL2.5-8B [9], and LLaVA-Video-7B [53]. Each of these models is evaluated with two typical strategies: uniform sampling and vectorized retrieval, where a CLIP-based retriever selects the top-K relevant frames based on the user query. In addition to VLMs, we benchmark Ava against SOTA Video-RAG frameworks, including VideoTree[47], VideoAgent[45], DrVideo[29], and VCA[48]. Among these, VideoTree, VideoAgent, and VCA are built upon GPT-4o, while DrVideo leverages GPT-4.
@@
Figure 8: The accuracy achieved by Ava and the baselines across typical query categories on LVBench: Temporal Grounding (TG), Summarization (SU), Reasoning (RE), Entity Recognition (ER), Event Understanding (EU), and Key Information Retrieval (KIR).
7.3 Overall Evaluation
7.3.1 Overall Performance
Fig. 7 illustrates the overall accuracy achieved by Ava compared to various baselines on the LVBench, VideoMME-Long, and Ava-100 benchmarks. Across all three benchmarks, Ava consistently outperforms the baselines. Specifically, on LVBench, Ava delivers a remarkable 16.9% improvement, while on VideoMME-Long, it advances the SOTA by approximately 5.2%. On the Ava-100 benchmark, Ava achieves an accuracy of 75.8%, significantly surpassing all competing methods.
In detail, compared to video-RAG methods, Ava achieves improvements of 21% and 7.8% on LVBench and VideoMME-Long, respectively. When compared to vectorized retrieval-based methods, Ava demonstrates gains of 16.9% on LVBench and 20.8% on Ava-100. Furthermore, against uniform sampling baselines, Ava improves performance by approximately 19.6% and 26.9% on LVBench and Ava-100, respectively.
Notably, on Ava-100, when evaluated with extremely long videos, Ava maintains robust performance, whereas the baselines degrade significantly. This highlights the effectiveness of Ava in handling L4 video analytics tasks.
7.3.2 Performance on Different Query Categories
@@
We also evaluate the accuracy achieved by Ava across typical query categories on LVBench. As illustrated in Fig. 8, our approach achieves improvements of 16%, 5.3%, 35.6%, 21.2%, 17.5%, and 18.9% across six key task types: Temporal Grounding, Summarization, Reasoning, Entity Recognition, Event Understanding, and Key Information Retrieval, respectively, compared to the uniform sampling and vectorized retrieval baselines powered by Gemini-1.5-Pro. Notably, Ava demonstrates particularly strong performance on reasoning tasks, which require identifying causal relationships between events and linking preceding and succeeding events within the video. This highlights Ava’s ability to effectively locate and extract critical information from long videos, thereby enabling advanced L4 video analytics systems.
@@
Figure 9: The accuracy achieved by Ava and baselines across three benchmarks when utilizing different LLMs and VLMs.
7.3.3 Performance under Different Configurations
@@
Fig. 9 shows the performance of Ava using different models configurations for SA and CA. For SA, two models were used: Qwen2.5 14B and 32B. For CA, two models were used: Qwen2.5-VL-7B and Gemini-1.5-Pro. The results show that across the three benchmarks, Ava using Gemini-1.5-Pro for CA achieved improvements of 18.9%, 5.2%, and 20.8% respectively compared to the best baseline result using the same model, while using Qwen2.5-VL-7B yielded improvements of 13%, 7.2%, and 15% respectively, fully demonstrating the effectiveness of our method. Notably, even when only using Qwen2.5-32B and Qwen2.5-7B based on the textual content from EKG without accessing raw frames, Ava can surpass the performance of Qwen2.5-VL-7B on the three benchmarks and also outperform most models shown in Figs. 7(a), 7(b), and 7(c).
7.3.4 Performance on Different Video Lengths
@@
Figure 10: The accuracy achieved by Ava and the baselines across varying video lengths via concatenating videos from LVBench.
@@
To evaluate the robustness of Ava with respect to video length, we conducted experiments on videos of varying durations. Specifically, sequences of 3.3, 6.6, and 10 hours were created by concatenating videos from the VideoMME-Long benchmark. Performance was measured using identical questions across these varying video lengths. As illustrated in Fig. 10, both Qwen2.5-VL-7B and Gemini-1.5-Pro baselines exhibit significant performance degradation as video length increases. When extended to 10 hours, their performance declines by 4.6% and 8.2%, respectively, under the uniform sampling method, compared to the original VideoMME-Long benchmark. For the vectorized retrieval setting, the performance drops are 4.6% and 5.5%, respectively. These results highlight the limitations of these methods in scaling effectively with increasing video length. In contrast, Ava consistently maintains stable performance across all video lengths, underscoring its robustness and scalability in handling video data of any duration.
7.3.5 System Overhead